# Chirality induced phase separation in active circle swimmers

### v1

@@ III.2 Dynamic of the system
Figure 6: (color online) $(a-c)$ MSD $\langle\Delta r^{2}(t)\rangle$ vs $t$ plots for the parameters where we find mixing of two types of particles and $(d-f)$ show the snapshots for the same parameters as for $(a-c)$ respectively. The arrows in the snapshots represent the orientation of particles and colors have the same meaning as in Fig. 2(I). $(a,d),(b,e),$ and $(c,f)$ are for the set of parameters $(v=0.5,\Omega=0.01),(v=0.01,\Omega=0.05)$ and $(v=0.001,\Omega=0.5)$.
@@
In Fig. 6(a-c) we again show the snapshots of the system for three different combinations of activity and chirality $(v_{0},\Omega)=(0.5,0.01),(0.01,0.05)$ and $(0.001,0.5)$ respectively. The three cases are chosen as follows: (a) When the trajectory of the particle is a nearly straight line (Fig. 2(I-II)(d) 3(d)). This case shows mixed flocking behavior, as seen by the ballistic nature of $MSD$ in Fig. 6(a). However, in the end, the $MSD$ shows saturation due to finite chirality in the system. This clearly explains the enhanced global order parameter $\Phi$ and small $PSOP$ for this set of parameters. In Fig. 6(e), we show the snapshot for $v_{0}=0.01$ and $\Omega=0.05$, where chirality starts to dominate and particles start moving in a small radius circle. Since the $r_{0}<R$, the particles get a chance to mix and rotate in small clusters. The $MSD$ shows the trapped motion at late times due to the rotational dynamic as shown in Fig. 6(b). Now as we decrease $v_{0}=0.001$ and increase $\Omega=0.5$ as shown in the snapshot in Fig. 6(f), chirality is dominant and particles mainly rotate about their axis without interacting with their neighbors. Hence, the system remains in homogeneous mixed phases with a very small magnitude of $MSD$ in comparison to Fig. 6(b). The dynamics of the particle are very different for the two cases Figs. 6(b) and (c). In Fig. 6(b), when both activity and chirality have moderate values but activity dominates, particles can form mixed oriented clusters with a rotational tendency, leading to a very small $MSD$ exponent $\alpha<<1$. On the other hand, for Fig. 6(c), when activity is very small and chirality is dominating, particles mainly rotate about their own axis with small random displacement from their mean position, resulting in a small magnitude of $MSD$ and subdiffusive dynamics in the system
@@ III.3 Phase diagram
We have analyzed the steady-state features of the system with respect to activity and chirality, and we drew the phase diagram in the plane of activity $v_{0}$ and chirality $\Omega$ as shown in Fig. 7. The colors in each symbol represent the magnitude of $PSOP$ and the different symbols are for four different phases found in the system based on the relative activity and chirality. We have already discussed the detailed characteristics of the system with respect to these different phases. Now based on their characteristics, here we name them and summarise. For very small chirality ($\Omega\leq 0.01$) and across the entire range of activity, the particles exist in a mixed state with small correlation length and $PSOP$ and flock in a common direction with moderate to large values of $GOP$, we refer to this phase as the mixed flocking phase $(MFP)$ and shown by the circles in Fig. 7. When the activity is small ($v_{0}<0.001$) and there is finite chirality, the particles rotate about their axis and both types of particles remain mixed, resulting in small values of $GOP$, $PSOP$, and $\mathcal{L}$. Such phase is shown by the pentagons in Fig. 7 and we named this phase and homogeneous rotating phase $(HRP)$. We have observed interesting phases emerging when activity and chirality compete. When the ratio $r_{0}<1$, the two types of particles get phase separated in small clusters, and both $PSOP$ and $GOP$ remain at moderate values. We define this phase as the microscopic phase separated phase $(MiPS)$, marked by triangles in Fig. 7. On the other hand, when the ratio is larger than $R$, macroscopic clusters of different types of particles are formed, and both $PSOP$ and $\mathcal{L}$ are large. This phase is named the macroscopic phase separated phase $(MaPS)$ and is represented by diamonds in Fig. 7. In the phase diagram, we have also shown the line with $r_{0}=R=1$. We can see that a simple mean field argument of the radius of the trajectory of particles can mimic the macroscopic phase separation in the system. Except for very small $\Omega=0.01$ and $v_{0}$.
@@
Figure 7: (color online) Phase Diagram for different phases in $(\Omega,v_{0})$ plane. The circle shows the mixed flocking $(MFP)$, the pentagon shows the homogeneous rotating $(HRP)$, the triangle shows the microscopic phase separated $(MiPS)$, and the diamond shows the macroscopic phase separated $(MaPS)$. The color bar shows the magnitude of the $PSOP$ $\Lambda$. The black dotted line shows the $\frac{v_{0}}{\Omega}=1=R$ line.
Till now we have discussed the steady properties of different phases. The macroscopic phase separated rotating phase $(MaPS)$ is the most interesting phase, where the two types of particles start to rotate in different clusters. Now we ask the question, what is the kinetics of the $(MaPS)$, when we start from the initial random disordered and mixed configuration?
@@
Figure 8: (color online) Shows the time-dependent density difference $\Delta\rho$ snapshots for different $v_{0}$ and $\Omega$ as $(v_{0},\Omega)=(0.1,0.06),(0.3,0.10),(0.5,0.17)$ for $t=2^{2},2^{6},2^{10}$ and $2^{16}$ and colour palette is same as in Fig.2(II).
@@
Figure 9: $(a)$ $v_{0}=0.1$, $(b)$ $v_{0}=0.2$, $(c)$ $v_{0}=0.3$ and $(d)$ $v_{0}=0.4$ and $(e)$ $v_{0}=0.5$ is two-point correlation function $C(r,t)$ vs. $r$ at different times. Insets: the scaled $C(r,t)$ vs scaled distance $r/\mathcal{L}$ for all five cases. $(f)$ Plot of correlation length $\mathcal{L}$ vs. $t$ in the growing stage on $\log-\log$ scale for different $(v_{0},\Omega)$ values.
@@
Figure 10: (a) Cartoon of the trajectories of two particles for the asymmetric combinations of chirality. The color code and keys are the same as for Fig. 4. The snapshots on the right side of each panel show a part of the system for the same set of parameters. The keys of the snapshots are the same as for Fig. 6. $(a-c)$ is for the set of parameters $(v_{0},\Omega^{+},\Omega^{-})=(0.5,0.1,-0.3),(0.5,0.1,0)$ and $(0.3,0.1,0)$ respectively.
III.4 Kinetics of $MaPS$ phase
@@
To understand the kinetics, we first observe the system’s time evolution. In Fig.8 we show the snapshot of the local density difference $\Delta\rho({\bf r})$ at four different times $t=2^{2}$, $2^{6}$, $2^{10}$ and $2^{16}$ for $(v_{0},\Omega)=((0.1,0.06),(0.3,0.10),(0.5,0.17))$. The choice of parameters is only those combinations of activity and chirality where the $PSOP$ and $\mathcal{L}$ shows the maxima in Fig. 5(a-b). With time the phase separation among the two types of particles increases as can be seen by the appearance of bright spots in the last two columns of Fig. 8. Also, the size of colored regions increases with time.
@@
Further, we show the plot of the two-point correlation function of local density difference as defined previously in subsection III.1. In Fig. 8(a-d) we show the plot of $C(r,t)$ for four different combinations of $(v_{0},\Omega)$ the same for which the snapshots are shown in Fig. 8. With time correlation increases and further we calculated the correlation length $\mathcal{L}(v_{0},\Omega,t)$. In the insets of Fig. 9(a-e) we show the scaled correlation function with scaled distance $r/\mathcal{L}$ for five different times $t=2^{2},2^{4},2^{6},2^{8},2^{10}$. The collapse of data for the scaled plots will represent the dynamic scaling in the system. Fig. 9(f) shows the variation of correlation length $\mathcal{L}(v_{0},\Omega,t)$ vs. time $t$ for different combinations of $(v_{0},\Omega)$. For all the cases, initially, $\mathcal{L}$ grows with time with a power $t^{\gamma}$, with $\gamma\sim 0.25$, and at late times the growth of $\mathcal{L}(v_{0},\Omega,t)$ slows as the system approaches the steady state.
III.5 Assymetric mixture
@@
Till now our focus was on the symmetric mixture of two types of chiral particles. Now we look for a case, where the magnitude of chirality of two types of particle, differ from each other. Hence we call it asymmetric mixture. In this case, again $50\%$ of the particles exhibit clockwise $(CW)$ chirality, while the remaining $50\%$ exhibit anticlockwise $(ACW)$ chirality, but the particles’ chirality magnitudes are not equal, indicating that the orientation of the two types of particles is not exactly opposite to each other. We will discuss the three combinations of chirality and activity. (i) As shown in Fig. 10, $ACW$-chiral particles ($\Omega^{-}=0.3$) exhibit higher chirality compared to $CW$-chiral particles ($\Omega^{+}=0.1$). In this case, chiral particle trajectories are a mixture of Case-II and Case-III as shown in Fig.4(b, c), and previously discussed in subsection III.1. Due to their smaller curvature, blue ($ACW$) particles spend more time in the interaction range than red ($CW$) particles with larger curvature.
Consequently, particles with $-ve$ chirality form one tight cluster and rotate due to their chirality, while $+ve$ chiral particles rotate in a dispersed manner in the system, as shown in the snapshot in Fig.10(a). The two other cases are (ii) when the one type particle is achiral $\Omega^{-}=0$ and $\Omega^{+}=0.1$ and $v_{0}=0.3$ for both types of particles. In Fig.10(b) we show the trajectory of two particles and the system snapshot at some fixed time in the steady state. In this scenario when activity is $v_{0}=0.3$, both types of particles form the mixed cluster and the mixture shows the chiral behavior. This happens because if the hopping step of particles is smaller than the $R$ then particles have a higher tendency to remain within the interaction range for some time and get a chance to interact with the mixture. Also curvature of chiral particles is larger resulting in them following the circular trajectory of the chiral particles, and the non-chiral (blue) particles rotate due to the chirality of the $+ve$ chiral particles as shown in the snapshot of Fig.10(b). However for the third case (c) when $v_{0}=0.5$, now the step size of particles is increased and this leads to the non-chiral particles leaving the surroundings of chiral particles very soon and they escape from the cluster of chiral particles as shown in Fig. 10(c).
@@ IV Discussion
We comprehensively studied a mixture of two types of chiral active particles. Most of our focus was on the symmetric mixture, where the two types of particles have the same magnitude of chirality but different signs. One type of particle rotates clockwise ($CW$) and the other type rotates counterclockwise ($ACW$). We used the activity $v_{0}$ and the magnitude of chirality $\Omega$ as the two control parameters in our model. The density of both types of particles was kept the same in the system. The magnitude $\Omega$ measures the rotational nature of particles as well as the difference between the two types of particles. For small $\Omega$, the difference in chirality between the two types of particles is small, causing them to remain in a mixed phase and move together in the same direction, Hence the system is found to be a mixed flocking phase $(MFP)$. For small activity and finite chirality, chirality dominates and the particles remain mixed, rotating about their axis without much translational motion. This phase is named the homogeneous rotating phase $(HRP)$. Interestingly, the system exhibits macroscopic phase separation and the rotating phase when activity and chirality are comparable. The phase separation in the system can be explained by defining the ratio $r_{0}=\frac{v_{0}}{\Omega}$, such that when the ratio is larger than the interaction range, particles get phase separated in macroscopic clusters. Each cluster rotates globally in one direction. The chirality of the majority of particles in the cluster decides the direction of rotation. Further when the ratio $r_{0}$ is smaller than the interaction range, then the two types of particles get phase separated in microscopic clusters. The clusters have the same tendency as found in macroscopic rotating clusters. The above two phases are named macroscopic phase separated $(MaPS)$ and microscopic phase separated $(MiPS)$ respectively. We also studied the kinetics of phase separation for the $MaPS$. We observed that the correlation length grows algebraically with time, with a growth exponent similar to the usual growth law for active systems Pattanayak et al. (2021); Redner et al. (2013), and the system exhibits good dynamic scaling for large activities $v_{0}$. Although the system shows the four different phases concerning activity and chirality and interesting dynamic patterns, we discussed the explanation of the four phases in detail with the help of simple mean field understanding using the radius of the particle trajectory and the interaction radius.
Furthermore, we explored asymmetric mixtures for a few combinations of activity and chirality, where the symmetric mixture system exhibits macroscopic phase-separated rotating clusters. We considered two cases where the chirality of one type of particle is smaller than the other, and two different activities are chosen, $v_{0}=0.3$ and $0.5$. We found two different phases: mixed rotating and rotating flocking phase, respectively.
@@
Our study provides a deep understanding of different phases for the mixture of chiral particles. The different phases we observed here can be useful to understand the patterns and dynamics observed in chiral active particles. The present study is limited to the point particles. It is interesting to explore the behavior of the system, where particles have finite size and shape which will make it closer to the many natural systems.
Acknowledgement

### v2

@@ III.2 Dynamics of the system
Figure 7: $(a,d,g)$ MSAD $\langle\Delta\theta^{2}(t)\rangle$ $vs.$ $t$ and $(b,e,h)$ MSD $\langle\Delta r^{2}(t)\rangle$ $vs.$ $t$ plots for the parameters where we find mixing of two types of particles and $(c,f,i)$ show the snapshots for the same parameters. The arrows in the snapshots represent the orientation of particles and colors have the same meaning as in Figs.2I (a,b,c)(a,b,c),$(d,e,f),$ and $(g,h,i)$ are for the set of parameters $(v=0.5,\Omega=0.01),(v=0.01,\Omega=0.05)$ and $(v=0.001,\Omega=0.5)$.
@@
We define the Mean Square displacement $(MSD)$ and Mean square angular displacement $(MSAD)$ as $\langle\Delta r^{2}\rangle(t)=\dfrac{1}{N}\langle\sum_{i=1}^{N}|\textbf{r}_{i}(t+t_{0})-\textbf{r}_{i}(t_{0})|^{2}\rangle$ and $\langle\Delta\theta^{2}\rangle(t)=\dfrac{1}{N}\langle\sum_{i=1}^{N}|\theta_{i}(t+t_{0})-\theta_{i}(t_{0})|^{2}\rangle$ respectively, where $\textbf{r}_{i}(t)$ and $\theta_{i}(t)$ are the position and angle of the $i^{th}$ particle at time $t$, $t_{0}$ represents the reference time of the measurement. Here, $\langle..$$\rangle$ indicates averaging over different values of $t_{0}$ in the steady-state trajectory. Due to periodic boundary conditions in both directions, the motion happens without any confinement, and $MSD$ has no upper bound. Hence, the behavior of $MSD$ and $MSAD$ with respect to time will show the translational and angular dynamics of the particles for different parameters. Generally, $MSD$ and $MSAD$ follow a power-law dependence on $t$. $\langle\Delta r^{2}\rangle\sim t^{\alpha}$ and $\langle\Delta\theta^{2}\rangle\sim t^{\beta}$, where the exponent $\alpha=1$ corresponds to diffusive motion, $\alpha<1$ to subdiffusive motion, and $\alpha>1$ to superdiffusive motion, while for ballistic motion, one has $\alpha=2$. Similarly, $\beta=1$ represents the random reorientation, and $\beta=2$ shows the persistent rotation of particles. In Figs.7(c,f,i), we again show the snapshots of the system for three different combinations of activity and chirality $(v_{0},\Omega)=(0.5,0.01),(0.01,0.05)$ and $(0.001,0.5)$ respectively. The three cases are chosen as follows: (a) When the trajectory of the particle is a nearly straight line (Figs.2(I-II)(d) 3(d)). This case shows mixed flocking behavior, as seen by the ballistic translational motion through $MSD$ and random reorientation through $MSAD$, as shown in Figs.7(a-c). However, at late times, the $MSD$ shows saturation due to finite chirality in the system. This clearly explains the enhanced global order parameter $\Phi$ and small $PSOP$ for this set of parameters. In Fig.7(f), we show the snapshot for $v_{0}=0.01$ and $\Omega=0.05$, where chirality starts to dominate, and particles start moving in a small radius circle shown by $MSAD$ ballistic motion Fig.7(d). Since the $r_{0}<R$, the particles can mix and rotate in small clusters. The $MSD$ shows the trapped motion at late times due to the rotational dynamics, as shown in Fig.7(e). Now, as we decrease $v_{0}=0.001$ and increase $\Omega=0.5$ as shown in the snapshot in Fig.7(i), chirality is dominant, and particles mainly rotate about their axis as shown by the ballistic behavior of $MSAD$ Fig.7(g) without interacting with their neighbors. Hence, the system remains in a homogeneous phase with a minimal magnitude of $MSD$ compared to Fig.7(h). The dynamics of the particle are very different for the two cases Figs.7(e) and (f). In Fig.7(e), when both activity and chirality have moderate values, but activity dominates, particles can form mixed oriented clusters with a rotational tendency, leading to a very small $MSD$ exponent $\alpha<<1$. On the other hand, for Fig.7(h), when activity is minimal and chirality is dominating, particles mainly rotate about their axis with small random displacement from their mean position, resulting in a small magnitude of $MSD$ and subdiffusive dynamics in the system.
@@ III.3 Phase diagram
Figure 8: (color online)Phase Diagram for different phases in $(\Omega,v_{0})$ plane. The circle shows the mixed flocking $(MFP)$, the pentagon shows the homogeneous rotating $(HRP)$, and the diamond shows the condensed phase $(CP)$. The color bar shows the magnitude of the PSOPPSOP $\Lambda$. The black dotted line shows the $\frac{v_{0}}{\Omega}=1=R$ line
We have analyzed the steady-state features of the system concerning activity and chirality, and we drew the phase diagram in the plane of activity $v_{0}$ and chirality $\Omega$ as shown in Fig.8. The colors in each symbol represent the magnitude of $PSOP$, and the different symbols represent three phases of the system based on relative activity and chirality. We have already discussed the detailed characteristics of the system for these different phases. Now, based on their characteristics, we name and summarize them here. For very small chirality ($\Omega\leq 0.01$) and across the entire range of activity, the particles exist in a mixed state with small correlation length and $PSOP$ and flock in a common direction with moderate to large values of $GOP$. We refer to this phase as the mixed flocking phase $(MFP)$ as shown by the circles in Fig.8. When the activity is small ($v_{0}<0.001$) and there is finite chirality, the particles rotate about their axis, and both types of particles remain mixed, resulting in small values of $GOP$, PSOPPSOP, and $\mathcal{L}$. The pentagons show such a phase in Fig.8, and we named this phase and homogeneous rotating phase $(HRP)$. We have observed interesting phases emerging when activity and chirality compete, where particles undergo segregation and exhibit condensation rather than conventional phase separation. Within this condensed phase $(CP)$, the nature of condensation depends on the ratio $v_{0}/\Omega$. When $r_{0}>1$, particles maintain alignment over longer distances before their trajectories curve significantly, leading to dense condensation, where clusters absorb a substantial fraction of particles and exhibit strong localization. In contrast, when $r_{0}<1$, the strong rotational bias restricts long-range alignment, resulting in weak condensation, where clusters form but remain more loosely structured and dynamic. As shown in Fig.8 $PSOP$.
Till now, we have discussed the steady properties of different phases.
@@
Figure 9: (color online) Shows the time-dependent density difference $\Delta\rho$ snapshots for different $(v_{0}=0.1,0.3,0.5)$ and $\Omega=1$ for $t=2^{2},2^{6},2^{10}$ and $2^{14}$ and color palette is same as in Fig.2(II).
@@
Figure 10: (color online) $(a)$ $v_{0}=0.3$ and $(b)$ $v_{0}=0.6$, is two-point correlation function $C(r,t)$ $vs.$ $r$ at different times for chirality $\Omega=0.1$. Insets: the scaled $C(r,t)$ $vs.$ scaled distance $r/\mathcal{L}$ for all five cases.
Figure 11: (color online) Plot of correlation length $\mathcal{L}$ $vs.$ $t$ in the growing stage on $\log$-$\log$ scale for different $v_{0}$ at $\Omega=0.1$ values.
@@
Figure 12: (a) Cartoon of the trajectories of two particles for the asymmetric combinations of chirality. The color code and keys are the same as for Fig.4. The snapshots on the right side of each panel show a part of the system for the same set of parameters. The keys of the snapshots are the same as for Figs.6.(a(a-$c)$ is for the set of parameters $(v_{0},\Omega^{+},\Omega^{-})=(0.5,0.1,-0.3),(0.5,0.1,0)$ and $(0.3,0.1,0)$ respectively.
III.4 Kinetics of condensing clusters
@@
To understand the kinetics, we first observe the system’s time evolution. In Fig.9 we show the snapshot of the local density difference $\Delta\rho({\bf r})$ at four different times $t=2^{2}$, $2^{6}$, $2^{10}$ and $2^{14}$ for $(v_{0})=((0.1),(0.3),(0.5))$ and $\Omega=1$. The choice of parameters is only the combinations of activity and chirality where we get $(CP)$. With time, the phase segregation among the two types of particles increases, as can be seen by the appearance of bright spots in the last two columns of Fig.9. Also, the size of colored regions increases with time.
@@
Furthermore, we show the plot of the two-point correlation function of the local density difference as defined previously in Subsection III.1. In Figs.10(a,b), we show the plot of $C(r,t)$ for two different $(v_{0}=0.3,0.6)$. With time, the correlation increases, and we calculate the correlation length $\mathcal{L}(v_{0},\Omega,t)$. In the insets of Figs.10(a,b), we show the scaled correlation function with scaled distance $r/\mathcal{L}$ for different twelve times between $t=50$ to $400$. The collapse of data for the scaled plots will represent the dynamic scaling in the system. Fig.11 shows the variation in the length of the correlation $\mathcal{L}(v_{0},\Omega,t)$ versus time $t$ for different combinations of $(v_{0}=0.05,0.1,0.2,0.3,0.4,0.5,0.6$ and $0.7)$ at $\Omega=0.1$. For all cases, initially, $\mathcal{L}$ grows with time with a power $t^{\gamma}$, with $\gamma\sim 0.25$, and at late times the growth of $\mathcal{L}(v_{0},\Omega,t)$ slows as the system approaches a steady state.
III.5 Asymmetric mixture
@@
Till now, our focus has been on the symmetric mixture of two types of chiral particles. Now, we look for a case where the magnitude of the chirality of two particle types differs. Hence, we call it an asymmetric mixture. In this case, again $50\%$ of the particles exhibit clockwise $(CW)$ chirality, while the remaining $50\%$ exhibit anticlockwise $(ACW)$ chirality, but the particles’ chirality magnitudes are not equal, indicating that the orientation of the two types of particles is not exactly opposite to each other. We will discuss the three combinations of chirality and activity. (i) As shown in Fig.12, $ACW$-chiral particles ($\Omega^{-}=0.3$) exhibit higher chirality compared to $CW$-chiral particles ($\Omega^{+}=0.1$). In this case, chiral particle trajectories are a mixture of Case-II and Case-III as shown in Figs.4(b, c), and previously discussed in subsection III.1. Due to their smaller curvature, blue $(ACW)$ particles spend more time in the interaction range than red $(CW)$ particles with larger curvature.
Figure 13: (color online) $\Lambda$ $vs.$ $|\Omega^{+}-\Omega^{-}|$ for different values of $v_{0}=\Omega^{+}$ (where, $|\Omega^{+}-\Omega^{-}|=|\Omega^{+}|+|\Omega^{-}|$). Each curve corresponds to a system where both the self-propulsion speed and the chirality of one species are fixed and equal. In contrast, the chirality of the second species $\Omega^{-}$ is varied to introduce asymmetry.
Consequently, particles with negative chirality form one strong cluster and rotate due to their chirality. In contrast, positive chiral particles rotate in a dispersed manner in the system, as shown in the snapshot in Fig.12(a). The other two cases are (ii) when the one-type particle is achiral $\Omega^{-}=0$ and $\Omega^{+}=0.1$ and $v_{0}=0.3$ for both types of particles. In Fig.12(b), we show the trajectory of two particles and the snapshot of the system at some fixed time in steady state. In this scenario, when activity is $v_{0}=0.3$, both particles form the mixed cluster, and the mixture shows chiral behavior. This happens because if the hopping step of particles is smaller than the $R$, then particles have a higher tendency to remain within the interaction range for some time and get a chance to interact with the mixture. Also, the curvature of the chiral particles is larger, resulting in them following the circular trajectory of the chiral particles, and the non-chiral particles (blue) rotate due to the chirality of the positive chiral particles, as shown in the snapshot of Fig.12(b). However, for the third case (c), when $v_{0}=0.5$, the increased step size of particles causes non-chiral particles to quickly leave the vicinity of chiral particles, as shown in Fig.12(c). To quantify the phase segregation among two types of particles for the asymmetric case, in Fig.13 we show the plot of phase separation order parameter $\Lambda$ vsvs difference in the chirality of two types of particles $|\Omega^{+}-\Omega^{-}|$. The $|\Omega^{+}|$ is kept fixed to $v_{0}$ and $v_{0}$ is varied between (0.20.5)(0.2-0.5). Hence, for one type of particle trajectory, the radius is roughly the same as the interaction radius. The chirality of the second type particles $\Omega^{-}$ varies from small to large values. When $|\Omega^{-}|<|\Omega^{+}|$, then both types of particles phase segregate and the system configuration is similar to that shown in Fig.12(a). The two types of particles phase separate into dense and dilute clusters. As we increase $|\Omega^{-}|$, the second type of particles starts to move in smaller circles, and the $\Lambda$ shows a dip. At larger $|\Omega^{-}|$, the second type of particles almost show confined motion to their location, and the system shows the partial phase segregation with $\Lambda~0.5$ value. The activity $v_{0}$ also plays a crucial role in the phase segregation. The moderate activity values like $v_{0}=0.2$ and $0.3$ revealed that increasing chirality asymmetry leads to a steady decrease in $\Lambda$. This occurs due to differences in rotational dynamics between particle types, which disrupt their alignment. At $v_{0}=0.2$, the system transitions through Cases II and III, where particles initially segregate due to favorable interactions. However, as chirality mismatch increases, the system shifts to Case IV, where particles primarily rotate in place, diminishing their ability to align and significantly reducing $\Lambda$. As activity increases further, particularly at $v_{0}=0.4$ and $0.5$, most of the system remains in the Cases II and III regimes. Enhanced velocities lead to larger effective trajectories, allowing for greater movement before deflections occur. This increased movement boosts the chances of particles aligning with similar neighbors, promoting phase segregation. Consequently, even with asymmetric chirality, the $\Lambda$ remains significant at higher velocities.
@@ IV Discussion
We comprehensively studied a mixture of two types of chiral active particles. Most of our focus was on the symmetric mixture, where the two types of particles have the same magnitude of chirality but different signs. One type of particle rotates clockwise $(CW)$, and the other rotates anticlockwise $(ACW)$. We used activity $v_{0}$ and magnitude of chirality $\Omega$ as two control parameters in our model. The density of both types of particles was kept the same in the system. The magnitude $\Omega$ measures the rotational nature of particles and the difference between the two types of particles. For small $\Omega$, the difference in chirality between the two types of particles is slight, causing them to remain in a mixed phase and move together in the same direction. Hence, the system is a mixed flocking phase $(MFP)$. For small activity and finite chirality, chirality dominates, and the particles remain mixed, rotating about their axes without much translational motion. This phase is named the homogeneous rotating phase $(HRP)$. Interestingly, the system exhibits a condensed phase when activity and chirality are comparable. When $r_{0}>1$, particles maintain alignment over longer distances before their trajectories curve significantly, leading to dense condensation, where clusters absorb a substantial fraction of particles and exhibit strong localization. In contrast, when $r_{0}<1$, the strong rotational bias restricts long-range alignment, resulting in weak condensation, where clusters form but remain more loosely structured and dynamic.
We also studied the kinetics of phase segregation for the condensation clusters. We observed that the correlation length grows algebraically with time, with a growth exponent similar to the usual growth law for active systems Pattanayak et al. (2021); Redner et al. (2013), and the system exhibits good dynamic scaling. Although the system shows the three phases concerning activity and chirality and interesting dynamic patterns, we discussed the explanation of the three phases in detail with the help of a simple mean-field understanding using the radius of the particle trajectory and the interaction radius.
Recent studies Ai et al. (2023, 2018); Li et al. (2024) have shown that phase segregation can arise in binary mixtures of chiral active particles with explicit excluded-volume interactions. In such systems, steric repulsion plays a crucial role in driving motility-induced phase separation, even in the absence of chirality or alignment. In contrast, our model considers point-like particles without volume exclusion, where dynamic clustering and phase segregation emerge purely from the interplay of velocity alignment and opposite chirality. This allows us to isolate chirality-induced effects and characterize the resulting condensed phases as a consequence of curvature-driven self-organization, rather than steric self-trapping. Furthermore, we explored asymmetric mixtures for a few combinations of activity and chirality, in contrast to the symmetric mixtures, where the system exhibits condensed phase segregation and rotating clusters. We considered two cases where the chirality of one type of particle is smaller than the other, and two different activities are chosen, $v_{0}=0.3$ and $0.5$. We found two phases: mixed rotating and rotating flocking phases.
@@
Our study provides a deep understanding of the different phases of the mixture of chiral particles. The various phases we observed here can help us understand the patterns and dynamics observed in chiral active particles. The present study is limited to the point particles with a uniform magnitude of chirality. It would be interesting to study the system with inhomogeneous chirality.
Author Contribution
The first author, DK, conducted the simulations, wrote the code, performed the result analysis, and generated the data. DK also wrote the initial draft of the manuscript. SM designed the model, assisted with coding, contributed to the discussion of the results, and prepared the final version of the manuscript.
Conflict of interest
There are no conflicts of interest to declare.
Acknowledgment